# Searching for GEMS: The Occurrence of Giant Planets orbiting M-dwarfs within 100 pc

### v1

@@ 3.1 Preparation of TGLC Light Curves
We use TGLCs (Han & Brandt, 2023a) for our analysis in place of light curves from the QLP (Huang et al., 2020) or TESS-SPOC (Caldwell et al., 2020). TESS-SPOC targets are selected with apparent $T$-mag $m_{\text{T}}<12$, with some extension to 13 depending on mission priorities (Sullivan et al., 2015; Stassun et al., 2018). Meanwhile QLP targets are selected to have $m_{\text{T}}<16-17$ (Huang et al., 2020). TGLC light curves exist out to $m_{\text{T}}\approx 18$, allowing our sample fuller coverage across faint M-dwarfs. Moreover, while the eleanor package (Feinstein et al., 2019) provides access to light curves for stars not covered by QLP or TESS-SPOC, it does not correct for dilution in the photometry due to nearby stars. This omission leads to significantly underestimated transit depths and biases in planet radius estimates for crowded fields. Previous work has documented these issues: for example, Kanodia et al. (2023b) and Kanodia et al. (2024) demonstrate that eleanor-based fits can mischaracterize transits in the presence of nearby stars. Meanwhile the TGLC pipeline, styled tglc, deblends crowding effects and provides complete sector-by-sector coverage. It has also demonstrated statistically accurate transit depths after correcting dilutions, making it well suited for robust occurrence rate studies across the TESS field (Han et al., 2025).
We retrieved 477,999 TGLCs from the TESS Primary Mission (PM) and First Extended Mission (EM1), covering all 149,316 stars in our sample. [^7] The processing of the TESS second extended mission (EM2) with tglc is ongoing. At the time of writing, the light curves analyzed in this study through Sector 51 are available on MAST at https://archive.stsci.edu/hlsp/tglc. We generated these light curves using the tglc package, applying the same processing parameters as the publicly available TGLC products hosted by the Mikulski Archive for Space Telescopes High-Level Science Products (Han & Brandt, 2023b). Each light curve was extracted from a $150\times 150$ pixel full-frame image (FFI) cutout. The substantial $150\times 150$ pixel size provides a robust estimate of the TESS point spread function (PSF). At each epoch, tglc fitted a PSF model based on stellar positions and magnitudes from Gaia Data Release 3 (Gaia Collaboration et al., 2023b). tglc then decontaminated the FFI cutouts by subtracting a simulated image constructed solely from field stars and background flux located within the FFI cutout. For our analysis, we opted not to use the calibrated aperture flux (cal_aper_flux) in order to perform our own independent detrending steps. We instead used the aperture flux (aperture_flux), computed by summing and normalizing the flux of the central $3\times 3$ pixels of the decontaminated FFI cut at each timestamp.
@@
    Initial Masking and Smoothing: We began by generating a smoothed flux curve using a Savitzky-Golay filter (Savitzky & Golay, 1964) with a window length of 13 cadences and a cubic polynomial.
@@ 3.2.1 Primary Sinusoidal Signal Search (GLS)
First, we identified the strongest sinusoidal signal in the sigma-clipped light curve using a generalized Lomb-Scargle (GLS) periodogram (Zechmeister & Kürster, 2009). To mitigate the influence of common TESS systematics (e.g., momentum dumps), we performed this search after outlier masking, which reduces sensitivity to low-frequency instrumental trends. We refer to the period corresponding to the highest-significance peak as $P_{\text{GLS, 1}}$.
@@ 3.2.2 Primary Detrending
We then detrended the light curve using wotan (Hippke et al., 2019), with the aim of removing the sinusoidal signal detected above in full or in part. We employed the built-in ‘cosine’ method with a window length of 5 days (or 2.5 days for the aforementioned active targets). This step removed lower-frequency trends while preserving astrophysical signals such as rapid rotation signatures, flares, and transits. The effect of this primary detrending is shown in Figure 2.
@@ 3.2.4 Primary Transit Search (BLS)
We searched for periodic, box-shaped transit signals in the single-detrended light curve using a Box-Least Squares (BLS) analysis (Kovács et al., 2002). We used the lightkurve.to_periodogram function (Lightkurve Collaboration et al., 2018) to compute the BLS power spectrum by phase-folding the light curve across trial periods ranging from 1.0 to 12.0 days. [^8] While our search for GEMS is restricted to orbital periods between 1.0 and 10.0 days, we extended the trial period range by two additional days to avoid edge effects and ensure robust detection near the upper boundary., using an oversampling factor of 7 to reduce the risk of missing potential transits. We defined the expected transit duration (the width of our test box) as 0.05 days, or 1.2 hours. This served only as an approximation, as we fit for the true transit duration later. This step yielded an initial estimate of the transit period ($P_{\text{BLS, 1}}$), the time of first transit ($T_{\text{init, 1}}$), and the approximate transit depth.
We also generated an “Anti-BLS” periodogram using the inverse light curve (1/Flux), effectively searching for periodic brightening events instead of periodic dips. A match between the strongest detected periods in both the standard BLS and the “Anti-BLS” periodograms (within 2% of the BLS period) indicated that the signals were most likely sinusoidal and hence likely due to rotational modulation. This has been demonstrated to be more effective than trying to match the GLS period with that obtained from the BLS periodogram and has been used successfully in previous studies (e.g., Gan et al., 2023). See Section 3.3 for the quality flag associated with this check.
@@ 3.2.5 Primary batman Fit
We generated a transit model with batman, taking the $P_{\text{BLS, 1}}$, $T_{\text{init, 1}}$, and primary BLS transit depth at maximum power as priors for optimization. We also set the prior values for impact parameter $b$ and $a/R_{*}$ as 0.3 and 15.0 respectively. We assumed an inclination angle of $90\degree$, a circular orbit ($e=0$, $\omega_{\star}=90^{\circ}$), and quadratic limb-darkening coefficients from Claret (2017). We fit the phase-folded data using scipy.optimize.minimize (Virtanen et al., 2020) to refine the coarse period estimate ($P_{\text{BLS, 1}}$), initial transit time estimate ($T_{\text{init, 1}}$), and transit duration estimate. We optimized the BLS fit with the Nelder-Mead method and a convergence tolerance of $10^{-6}$ (Nelder & Mead, 1965). The Nelder-Mead method is particularly useful for fitting complex models without the need for gradients, but has the disadvantage of being relatively computationally expensive (see Section 5.1).
@@ 3.2.6 Transit Masking and Secondary Detrending
We used the estimated transit parameters identified in the previous step to temporarily mask all detected transits in the single-detrended light curve. We then detrended the remaining data using a finer window length of 0.5 days (or 0.25 days if $P_{\text{GLS, 2}}<1.0~\mathrm{day}$). Using the numpy package (Harris et al., 2020), we interpolated the masked regions with a one-dimensional linear interpolation, which effectively bridges short data gaps like those caused by transits. We subtracted the resulting trend from the single-detrended data to produce a twice-flattened light curve. Figure 2 shows an example of this step. While the single-detrended light curve retains some high-frequency modulation, the secondary detrending removes it cleanly and yields a satisfactorily flat curve that preserves the transit signal.
@@ 3.3 Target Classification
  • DensityFlag: True if the fitted stellar density from Section 3.2.7 falls outside specified, customizable bounds. Assuming a circular orbit or with knowledge of orbital eccentricity $e$, a genuine planetary transit provides an opportunity to measure the host star’s density (Seager & Mallén-Ornelas, 2003; Winn, 2010). [^10] We chose to fit circular models due to the short orbital periods being probed here. In our full sample, the star possessed of the lowest density was TIC 307928780 ($3.12$ g cm-3), and the star with the highest density was TIC 165086573 ($92.59$ g cm-3). We adopted a lower limit of 1 $\textrm{g~cm}^{-3}$ and an upper limit of 200 $\textrm{g~cm}^{-3}$, with the limits conservatively defined to exclude only the poorest fits.
@@ 4.1 Stage 0: Initial TESS-miner Result Breakdown
Figure 3: The cascade of candidates from the initial sample through our final GEMS list. Steps in our search include analysis with TESS-miner (Stage 0; yellow), auto-vetting (Stage 1; orange), by-eye and otherwise supervised inspection steps (Stage 2; purple), and spectroscopic validation (Stage 3; blue). Final results are divided into false positives (FPs; white), previously confirmed GEMS (bright green), and new GEMS candidates (light green).
Figure 4: Map of all targets. Poor Fits are shown in pale turquoise, Good Fits in dark blue, Bad Data in red, and targets lacking TGLC data in black points. [^13] Note that legend markers are enlarged for clarity. Black points indicate stars lacking TESS data and should not be confused with the black polygons denoting the Kepler and Roman survey footprints overlaid on the figure. Spectroscopically vetted targets (Table 2) are overplotted: FPs as red triangles, known GEMS as dark blue astroids, and new candidate GEMS as yellow stars. The Kepler footprint (black bars; Mullally et al., 2016) and Nancy Roman Galactic Bulge Time Domain Survey footprint [^14] Adopted centers from https://github.com/mtpenny/gbtds_optimizer/blob/main/field_layouts/layout_40395.centers (black polygons indicated by arrow; Spergel et al. (2015); Penny et al. (2019); ROTAC (2025)) are shown for reference.
@@ 4.2 Stage 1: Auto-vetting Steps
    BLS Power Cut: We removed all objects from consideration whose maximum BLS power was below 100. We use BLS power here instead of traditional Gaussian S/N because it is specifically designed to detect periodic, box-shaped transits by phase-folding light curves. This makes it more sensitive and robust to real transit signals in noisy, irregular data. The formulation for BLS power is defined by Equation 11 in Kovács et al. (2002) as the square of the “effective S/N” $\alpha=\frac{\delta}{\sigma}\sqrt{nq}$, where $\delta$ is the transit depth, $\sigma$ is the per-point noise, $n$ is the total number of data points, and $q$ is the fractional transit duration.
@@
    V-Shaped Cut: We defined an additional cut on V-shaped transits using the same method as appears in the LEO-vetter auto-vetting package (Kunimoto, 2024). If $b+R_{p}/R_{\star}>1.5$ where $b$ is the fitted impact parameter, we classified the light curve as V-shaped and removed it from consideration. A similar metric was used in the Kepler mission (Thompson et al., 2018) to reliably identify EBs as too deep (large $R_{p}/R_{\star}$) or extremely grazing (large $b$). This threshold is the default choice for LEO-vetter, and preserved our test sample of all known GEMS within 200 pc. We found that 24.3% of surviving candidates had fitted impact parameters $b>0.85$ even after applying this cut, alleviating concerns that it might exclude grazing planets from our search.
@@ 4.3 Stage 2a: Manual Vetting
  • FP due to an eccentric EB: A handful of objects were eccentric EB systems, where the secondary transit signal did not occur at the exact midpoint between primary transits. TESS-miner does not yet account for this type of signal pattern. [^15] We note here that M-dwarf/M-dwarf binaries, of which these objects are prime examples, tend to have higher primary to secondary mass ratios, as well as lower eccentricities, than other EB spectral pairings (Fontanive et al., 2018; Dupuy & Liu, 2017). This makes eccentric EBs quite rare in our sample, as well as interesting targets for future research in that field.
  • FP due to contact binary: Contact binaries display constant flux modulation due to the distortion of the component stars (Mattei & Saladyga, 1999), and periods of less than 0.7 days (Hilditch, 2001). [^16] In these cases, TESS-miner identified an alias of the true period which was within the search bounds of 1.0 to 12.0 days. Their signal often includes equal primary and secondary transit depths due to roughly equivalent component radii (Kuiper, 1941; Lucy, 1967). These features combine to produce a characteristic scalloped pattern.
@@ 4.5 Stage 2c: Stellar Density Estimation with juliet
We used the Bayesian modeling tool juliet (Espinoza et al., 2019) to fit the double-detrended light curves from TESS-miner for the 35 candidates that passed pixel-level vetting. juliet employs dynamic nested sampling via the dynesty package (Speagle, 2020) to sample the full posterior distribution. Unlike the Nelder-Mead optimizer used for preliminary fits in TESS-miner, juliet supports informative priors and explicit modeling of systematics, enabling more robust discrimination between planetary transits and EBs.
For this work, we used the default dynesty convergence criterion of $\Delta\ln Z<0.01$. We generated transit models using batman (Kreidberg, 2015), adopting a quadratic limb-darkening law and sampling coefficients from uniform priors following Kipping (2013). To account for finite integration times, we applied a supersampling factor of 30. We placed Gaussian priors ($\sigma=0.1~\mathrm{days}$) on the period and mid-transit time derived from TESS-miner. All fits assumed circular orbits ($e=0$, $\omega_{\star}=90^{\circ}$) and a fixed dilution factor of unity. [^17] Although GEMS may retain spin-orbit misalignment and eccentricity (Section 1), only one confirmed GEM to date demonstrates appreciable eccentricity (TOI-6330 b; Hotnisky et al. (2025)), and it does not lie within 100 pc. We fixed the dilution factor because TESS photometry alone (i.e., without high-resolution imaging or multi-band photometry) cannot constrain contamination. The assumption of zero eccentricity yields smaller radius estimates and is a conservative choice when modeling low-precision TESS data (e.g., Kipping, 2008). The photometric model included a sector-specific white-noise jitter term added in quadrature to the reported uncertainties. We jointly fitted all available TESS sectors per target, allowing sector-specific baselines and jitter terms while sharing limb-darkening parameters. The final fits and phase-folded light curves appear in Figure 6.
Figure 6: juliet fits for all TICs referenced in Table 2 as a new candidate (row 1) or astrophysical FPs identified in the spectroscopic vetting stage (rows 2-5).

### v3

@@ 3.1 Preparation of TGLC Light Curves
We use TGLCs (Han2023) for our analysis in place of light curves from the QLP (huang_photometry_2020) or TESS-SPOC (caldwell_tess_2020). TESS-SPOC targets are selected with apparent $T$-mag $m_{\text{T}}<12$, with some extension to 13 depending on mission priorities (sullivan15; stassun18). Meanwhile QLP targets are selected to have $m_{\text{T}}<16-17$ (huang_photometry_2020). TGLC light curves exist out to $m_{\text{T}}\approx 18$, allowing our sample fuller coverage across faint M-dwarfs. Moreover, while the eleanor package (feinstein_eleanor_2019) provides access to light curves for stars not covered by QLP or TESS-SPOC, it does not correct for dilution in the photometry due to nearby stars. This omission leads to significantly underestimated transit depths and biases in planet radius estimates for crowded fields. Previous work has documented these issues: for example, kanodia_toi-5205b_2023 and kanodia_searching_2024-1 demonstrate that eleanor-based fits can mischaracterize transits in the presence of nearby stars. Meanwhile the TGLC pipeline, styled tglc, deblends crowding effects and provides complete sector-by-sector coverage. It has also demonstrated statistically accurate transit depths after correcting dilutions, making it well suited for robust occurrence rate studies across the TESS field (Han2025_larger).
We retrieved 477,999 TGLCs from the TESS Primary Mission (PM) and First Extended Mission (EM1), covering all 149,316 stars in our sample. [^7] The processing of the TESS second extended mission (EM2) with tglc is ongoing. At the time of writing, the light curves analyzed in this study through Sector 54 are available on MAST at https://archive.stsci.edu/hlsp/tglc. We generated these light curves using the tglc package, applying the same processing parameters as the publicly available TGLC products hosted by the Mikulski Archive for Space Telescopes High-Level Science Products (tglc_hlsp). Each light curve was extracted from a $150\times 150$ pixel full-frame image (FFI) cutout. The substantial $150\times 150$ pixel size provides a robust estimate of the TESS point spread function (PSF). At each epoch, tglc fitted a PSF model based on stellar positions and magnitudes from Gaia Data Release 3 (gaia_collaboration_gaia_2023). tglc then decontaminated the FFI cutouts by subtracting a simulated image constructed solely from field stars and background flux located within the FFI cutout. For our analysis, we opted not to use the calibrated aperture flux (cal_aper_flux) in order to perform our own independent detrending steps. We instead used the aperture flux (aperture_flux), computed by summing and normalizing the flux of the central $3\times 3$ pixels of the decontaminated FFI cut at each timestamp.
@@
    Initial Masking and Smoothing: We began by generating a smoothed flux curve using a Savitzky-Golay filter (savitzky1964smoothing) with a window length of 13 cadences and a cubic polynomial.
@@ 3.2.1 Primary Sinusoidal Signal Search (GLS)
First, we identified the strongest sinusoidal signal in the sigma-clipped light curve using a generalized Lomb-Scargle (GLS) periodogram (zechmeister_kurster_2009). To mitigate the influence of common TESS systematics (e.g., momentum dumps), we performed this search after outlier masking, which reduces sensitivity to low-frequency instrumental trends. We refer to the period corresponding to the highest-significance peak as $P_{\text{GLS, 1}}$.
@@ 3.2.2 Primary Detrending
We then detrended the light curve using wotan (Wotan), with the aim of removing the sinusoidal signal detected above in full or in part. We employed the built-in ‘cosine’ method with a window length of 5 days (or 2.5 days for the aforementioned active targets). This step removed lower-frequency trends while preserving astrophysical signals such as rapid rotation signatures, flares, and transits. The effect of this primary detrending is shown in Figure 2.
@@ 3.2.4 Primary Transit Search (BLS)
We searched for periodic, box-shaped transit signals in the single-detrended light curve using a Box-Least Squares (BLS) analysis (kovacs_box-fitting_2002). We used the lightkurve.to_periodogram function (lightkurve) to compute the BLS power spectrum by phase-folding the light curve across trial periods ranging from 1.0 to 12.0 days. [^8] While our search for GEMS is restricted to orbital periods between 1.0 and 10.0 days, we extended the trial period range by two additional days to avoid edge effects and ensure robust detection near the upper boundary., using an oversampling factor of 7 to reduce the risk of missing potential transits. We defined the expected transit duration (the width of our test box) as 0.05 days, or 1.2 hours. This served only as an approximation, as we fit for the true transit duration later. This step yielded an initial estimate of the transit period ($P_{\text{BLS, 1}}$), the time of first transit ($T_{\text{init, 1}}$), and the approximate transit depth.
We also generated an “Anti-BLS” periodogram using the inverse light curve (1/Flux), effectively searching for periodic brightening events instead of periodic dips. A match between the strongest detected periods in both the standard BLS and the “Anti-BLS” periodograms (within 2% of the BLS period) indicated that the signals were most likely sinusoidal and hence likely due to rotational modulation. This has been demonstrated to be more effective than trying to match the GLS period with that obtained from the BLS periodogram and has been used successfully in previous studies (e.g., gan_occurrence_2023). See Section 3.3 for the quality flag associated with this check.
@@ 3.2.5 Primary batman Fit
We generated a transit model with batman, taking the $P_{\text{BLS, 1}}$, $T_{\text{init, 1}}$, and primary BLS transit depth at maximum power as priors for optimization. We also set the prior values for impact parameter $b$ and $a/R_{*}$ as 0.3 and 15.0 respectively. We assumed an inclination angle of $90\degree$, a circular orbit ($e=0$, $\omega_{\star}=90^{\circ}$), and quadratic limb-darkening coefficients from claret_limb_2017. We fit the phase-folded data using scipy.optimize.minimize (scipy) to refine the coarse period estimate ($P_{\text{BLS, 1}}$), initial transit time estimate ($T_{\text{init, 1}}$), and transit duration estimate. We optimized the BLS fit with the Nelder-Mead method and a convergence tolerance of $10^{-6}$ (neldermead1965). The Nelder-Mead method is particularly useful for fitting complex models without the need for gradients, but has the disadvantage of being relatively computationally expensive (see Section 5.1).
@@ 3.2.6 Transit Masking and Secondary Detrending
We used the estimated transit parameters identified in the previous step to temporarily mask all detected transits in the single-detrended light curve. We then detrended the remaining data using a finer window length of 0.5 days (or 0.25 days if $P_{\text{GLS, 2}}<1.0~\mathrm{day}$). Using the numpy package (harris_array_2020), we interpolated the masked regions with a one-dimensional linear interpolation, which effectively bridges short data gaps like those caused by transits. We subtracted the resulting trend from the single-detrended data to produce a twice-flattened light curve. Figure 2 shows an example of this step. While the single-detrended light curve retains some high-frequency modulation, the secondary detrending removes it cleanly and yields a satisfactorily flat curve that preserves the transit signal.
@@ 3.3 Target Classification
  • DensityFlag: True if the fitted stellar density from Section 3.2.7 falls outside specified, customizable bounds. Assuming a circular orbit or with knowledge of orbital eccentricity $e$, a genuine planetary transit provides an opportunity to measure the host star’s density (seager_unique_2003; winn_exotransits_2010). [^10] We chose to fit circular models due to the short orbital periods being probed here. In our full sample, the star possessed of the lowest density was TIC 307928780 ($3.12$ g cm-3), and the star with the highest density was TIC 165086573 ($92.59$ g cm-3). We adopted a lower limit of 1 $\textrm{g~cm}^{-3}$ and an upper limit of 200 $\textrm{g~cm}^{-3}$, with the limits conservatively defined to exclude only the poorest fits.
@@ 4.1 Stage 0: Initial TESS-miner Result Breakdown
Figure 3: The cascade of candidates from the initial sample through our final GEMS list. Steps in our search include analysis with TESS-miner (Stage 0; yellow), auto-vetting (Stage 1; orange), by-eye and otherwise supervised inspection steps (Stage 2; purple), and spectroscopic validation (Stage 3; blue). Final results are divided into false positives (FPs; white) and previously confirmed GEMS (green).
Figure 4: Map of all targets. Poor Fits are shown in pale turquoise, Good Fits in dark blue, Bad Data in red, and targets lacking TGLC data in black points. [^13] Note that legend markers are enlarged for clarity. Black points indicate stars lacking TESS data and should not be confused with the black polygons denoting the Kepler and Roman survey footprints overlaid on the figure. Spectroscopically vetted targets (Table 2) are overplotted: FPs as red triangles and known GEMS as dark blue astroids. The FPs denoted here include those identified via spectroscopic followup by this work (13) and by previous works (3). The Kepler footprint (black bars; 2016ascl.soft01009M) and Nancy Roman Galactic Bulge Time Domain Survey footprint [^14] Adopted centers from https://github.com/mtpenny/gbtds_optimizer/blob/main/field_layouts/layout_40395.centers (black polygons indicated by arrow; Spergel15; Penny2019; rotac2025) are shown for reference.
@@ 4.2 Stage 1: Auto-vetting Steps
    BLS Power Cut: We removed all objects from consideration whose maximum BLS power was below 100. We use BLS power here instead of traditional Gaussian S/N because it is specifically designed to detect periodic, box-shaped transits by phase-folding light curves. This makes it more sensitive and robust to real transit signals in noisy, irregular data. The formulation for BLS power is defined by Equation 11 in kovacs_box-fitting_2002 as the square of the “effective S/N” $\alpha=\frac{\delta}{\sigma}\sqrt{nq}$, where $\delta$ is the transit depth, $\sigma$ is the per-point noise, $n$ is the total number of data points, and $q$ is the fractional transit duration.
@@
    V-Shaped Cut: We defined an additional cut on V-shaped transits using the same method as appears in the LEO-vetter auto-vetting package (Kunimoto_2024). If $b+R_{p}/R_{\star}>1.5$ where $b$ is the fitted impact parameter, we classified the light curve as V-shaped and removed it from consideration. A similar metric was used in the Kepler mission (kepdr25) to reliably identify EBs as too deep (large $R_{p}/R_{\star}$) or extremely grazing (large $b$). This threshold is the default choice for LEO-vetter, and preserved our test sample of all known GEMS within 200 pc. We found that 24.3% of surviving candidates had fitted impact parameters $b>0.85$ even after applying this cut, alleviating concerns that it might exclude grazing planets from our search.
@@ 4.3 Stage 2a: Manual Vetting
  • FP due to an eccentric EB: A handful of objects were eccentric EB systems, where the secondary transit signal did not occur at the exact midpoint between primary transits. TESS-miner does not yet account for this type of signal pattern. [^15] We note here that M-dwarf/M-dwarf binaries, of which these objects are prime examples, tend to have higher primary to secondary mass ratios, as well as lower eccentricities, than other EB spectral pairings (Fontanive2018; dupuy_individual_2017). This makes eccentric EBs quite rare in our sample, as well as interesting targets for future research in that field.
  • FP due to contact binary: Contact binaries display constant flux modulation due to the distortion of the component stars (Mattei1999), and periods of less than 0.7 days (Hilditch2001). [^16] In these cases, TESS-miner identified an alias of the true period which was within the search bounds of 1.0 to 12.0 days. Their signal often includes equal primary and secondary transit depths due to roughly equivalent component radii (Kuiper1941; Lucy1967). These features combine to produce a characteristic scalloped pattern.
@@ 4.5 Stage 2c: Stellar Density Estimation with juliet
We used the Bayesian modeling tool juliet (espinoza_juliet_2019) to fit the double-detrended light curves from TESS-miner for the 35 candidates that passed pixel-level vetting. juliet employs dynamic nested sampling via the dynesty package (speagle_dynesty_2020) to sample the full posterior distribution. Unlike the Nelder-Mead optimizer used for preliminary fits in TESS-miner, juliet supports informative priors and explicit modeling of systematics, enabling more robust discrimination between planetary transits and EBs.
For this work, we used the default dynesty convergence criterion of $\Delta\ln Z<0.01$. We generated transit models using batman (kreidberg_batman_2015), adopting a quadratic limb-darkening law and sampling coefficients from uniform priors following kipping_efficient_2013. To account for finite integration times, we applied a supersampling factor of 30. We placed Gaussian priors ($\sigma=0.1~\mathrm{days}$) on the period and mid-transit time derived from TESS-miner. All fits assumed circular orbits ($e=0$, $\omega_{\star}=90^{\circ}$) and a fixed dilution factor of unity. [^17] Although GEMS may retain spin-orbit misalignment and eccentricity (Section 1), only one confirmed GEM to date demonstrates appreciable eccentricity (TOI-6330 b; hotnisky_2025_searching), and it does not lie within 100 pc. We fixed the dilution factor because TESS photometry alone (i.e., without high-resolution imaging or multi-band photometry) cannot constrain contamination. The assumption of zero eccentricity yields smaller radius estimates and is a conservative choice when modeling low-precision TESS data (e.g., Kipping2008). The photometric model included a sector-specific white-noise jitter term added in quadrature to the reported uncertainties. We jointly fitted all available TESS sectors per target, allowing sector-specific baselines and jitter terms while sharing limb-darkening parameters. The final fits and phase-folded light curves appear in Figure 6.
Figure 6: juliet fits for the 13 TICs cited in Table 2 as new astrophysical false positives by this work.